# Non-local planelike minimizers and $\Gamma$-convergence of periodic energies to a local anisotropic perimeter

### v1

@@ 9. Proof of the $\Gamma$-$\liminf$ inequality
and $\lambda(A)\leqslant\liminf_{\varepsilon\to 0}\lambda_{\varepsilon}(A),\quad\mbox{for every set\penalty 10000\ $A$ such that\penalty 10000\ $\lambda(\partial A)=0$.}$
@@
Then, let $\overline{x}$ be a regular point of $E$, namely a point $\overline{x}\in\partial E$ with (n1)(n-1)-density 11 and such that the blow-up of $E$ around $\overline{x}$ converges to $\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}$, e.g.
@@
and $\lim_{r\to 0^{+}}r^{-n}\int_{B_{2r}(\overline{x})}|\chi_{E}-\chi_{\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}}|dx=0,$
where $\omega_{n-1}:=\mathcal{H}^{n-1}(B^{n-1}_{1})$. We consider the Radon-Nikodyn derivative of $\lambda$ with respect to the (n1)(n-1)-dimensional Hausdorff measure at $\overline{x}$
@@
(9.3) $\frac{d\lambda}{d\mathcal{H}^{n-1}}(\overline{x})\geqslant\phi(\nu_{E}(\overline{x})).$
At this stage, let us consider subsequences $\{\varepsilon_{j}\}_{j}$, $\{r_{j}\}_{j}$, and $\{\eta_{j}:=\varepsilon_{j}/r_{j}\}_{j}$ such that
(9.4) $\begin{split}&\lim_{j\to+\infty}\frac{\lambda_{\varepsilon_{j}}(B_{r_{j}}(\overline{x}))}{\omega_{n-1}r_{j}^{n-1}}=\frac{d\lambda}{d\mathcal{H}^{n-1}}(\overline{x})\\ &\lim_{j\to+\infty}r_{j}^{-n}\int_{B_{2r_{j}}(\overline{x})}|\chi_{E}-\chi_{\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}}|dx=0,\\ \mbox{and}\quad&\lim_{j\to+\infty}\eta_{j}=0.\end{split}$
In what follows, in order to lighten the notation, we will write $E_{j}:=E_{\varepsilon_{j}}$, $\lambda_{\varepsilon_{j}}:=\lambda_{j}$, and $\nu:=\nu_{E}(\overline{x})$. We now focus on proving (9.3), that now is
@@
$\begin{split}&\frac{\lambda_{j}(B_{r_{j}}(\overline{x}))}{r_{j}^{n-1}}=\iint_{B_{1}(\eta_{j}\{\overline{x}/\varepsilon_{j}\})\times\mathbb{R}^{n}}\chi_{E^{\prime\prime}_{j}}(x)\chi_{E^{\prime\prime c}_{j}}(y)K_{\eta_{j}}(x,y)\,dx\,dy+\int_{E^{\prime\prime}_{j}\cap\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}}g_{\eta_{j}}(x)dx\\ &\qquad=\iint_{B^{\prime\prime}_{j}\times\mathbb{R}^{n}}\chi_{E^{\prime\prime}_{j}}(x)\chi_{E^{\prime\prime c}_{j}}(y)K_{\eta_{j}}(x,y)\,dx\,dy+\int_{E^{\prime\prime}_{j}\cap\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}}g_{\eta_{j}}(x)dx,\end{split}$
where, to be consistent with our notation, we wrote $B_{j}^{\prime\prime}:=B_{1}(\eta_{j}\{\overline{x}/\varepsilon_{j}\})$. Moreover, observe that
$\begin{split}&\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}=\bigcup\left\{\frac{\varepsilon_{j}(k+Q)-\overline{x}}{r_{j}}+\eta_{j}\left\{\frac{\overline{x}}{\varepsilon_{j}}\right\}\mbox{s.t. }k\in\mathbb{Z}^{n},\ \varepsilon_{j}(k+Q)\subseteq B_{r_{j}}(\overline{x})\right\}\\ &\qquad=\bigcup\left\{\eta_{j}\left(k-\left\lfloor\frac{\overline{x}}{\varepsilon_{j}}\right\rfloor+Q\right)\mbox{s.t. }k\in\mathbb{Z}^{n},\ \eta_{j}\left(k-\left\lfloor\frac{\overline{x}}{\varepsilon_{j}}\right\rfloor+Q\right)\subseteq B^{\prime\prime}_{j}(\overline{x})\right\}\\ &\qquad=\mathcal{Q}(B^{\prime\prime}_{j})_{\eta_{j}}.\end{split}$
@@
(9.7) $\begin{split}&\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{E^{\prime\prime}_{j}}-\chi_{\mathcal{I}_{\nu}}|dx\\ \leqslant\;&\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{E^{\prime\prime}_{j}}-\chi_{\{(x-\eta_{j}\{\overline{x}/\varepsilon_{j}\})\cdot\nu>0\}}|dx+\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{\{(x-\eta_{j}\{\overline{x}/\varepsilon_{j}\})\cdot\nu>0\}}-\chi_{\mathcal{I}_{\nu}}|dx\\ =\;&0.\end{split}$
@@
(9.8) $F_{j}:=(\eta_{j}E_{\nu}\setminus\mathcal{Q}(B^{\prime\prime}_{j})_{1})\cup(E^{\prime\prime}_{j}\cap\mathcal{Q}(B^{\prime\prime}_{j})_{1}).$
@@
(9.9) $\mathscr{F}_{\eta_{j}}(F_{j},B^{\prime\prime}_{j})\geqslant\mathscr{F}_{\eta_{j}}(\eta_{j}E_{\nu},B^{\prime\prime}_{j})\to\mathcal{H}^{n-1}(\partial\mathcal{I}_{\nu}\cap B_{1})\phi(\nu)=\omega_{n-1}\phi(\nu),$
as $j\to+\infty$.
@@ 10. Continuity of the stable norm $\phi$
$\begin{split}\varepsilon^{1-n}\zeta(\varepsilon):=&\left|S^{q}_{M+\sqrt{n}}\cap S^{p}_{M+\sqrt{n}}\cap\mathcal{R}_{\theta}E^{q}\cap\left[C^{p}/\varepsilon\setminus\mathcal{Q}(Q^{p}/\varepsilon)_{1}\right]\right|\\ &+\left|S^{q}_{M+\sqrt{n}}\cap S^{p}_{M+\sqrt{n}}\cap E^{q}\cap\left[C^{q}/\varepsilon\setminus\mathcal{Q}(Q^{q}/\varepsilon)_{1}\right]\right|,\end{split}$
@@
$\begin{split}&\phi(q)-\phi(p)\\ \geqslant&\liminf_{\varepsilon\to 0}-\varepsilon^{n-1}\|g\|_{\infty}\left|[E_{q}\cap S^{q}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{q}]\Delta[\mathcal{R}_{\theta}E_{q}\cap S^{p}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{p}]\right|.\end{split}$
@@
$\psi(\theta):=\limsup_{\varepsilon\to 0}\varepsilon^{n-1}\left|[E_{q}\cap S^{q}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{q}]\Delta[\mathcal{R}_{\theta}E_{q}\cap S^{p}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{p}]\right|,$
@@
$\phi(q)-\phi(p)\geqslant-\|g\|_{\infty}\psi(\theta).$
@@
Interchanging the roles of $p$ and $q$, we obtain again that
$\begin{split}&\phi(p)-\phi(q)\geqslant\liminf_{\varepsilon\to 0}-\|g\|_{\infty}\left|[E_{q}\cap S^{p}_{M+\sqrt{n}}\cap C^{p}]\Delta[\mathcal{R}_{-\theta}\cap S^{q}_{M+\sqrt{n}}\cap C^{q}]\right|\\ &\qquad\qquad=-\|g\|_{\infty}|J\mathcal{R}_{\theta}|\psi(\theta)=-\|g\|_{\infty}\psi(\theta),\end{split}$
@@ 11. Proof of the $\Gamma$-$\limsup$ inequality
We point out that, in light of Lemma 7.1, one can equivalently consider energies $\mathscr{F}_{\varepsilon}((\cdot),\Omega)$, $\mathscr{E}_{\varepsilon}((\cdot),\Omega)$, or even $\mathscr{L}_{\varepsilon}((\cdot),\Omega):=\mathcal{L}_{K_{\varepsilon}}((\cdot)\cap\Omega,(\cdot)^{c}\cap\Omega)$ in Theorem 1.13. Indeed, since the rescaled planelike $\varepsilon E_{p}$ converges to the straight line $\mathcal{I}_{p}$, as $\varepsilon\to 0$, such limit crosses transversely any Lipschitz domain. Therefore, the energy contribution due to the interactions with the exterior of $\Omega$ in the limit.
@@ Appendix A Norms subject to short-range interactions
We point out a useful equivalence of norms when a kernel is restricted to close-by interactions (see e.g. [MR5007887, Lemma 2.1] for related results). For the rest of this section, let $s\in(0,1/2)$.
@@ Appendix C Behavior of $\operatorname{P}_{K}$ in large domains
First, let us consider the space of Lipschitz function on $B_{\rho}^{n-1}$, for some $\rho>0$,
$\mbox{Lip}(B_{\rho}^{n-1}):=\{f:B_{\rho}^{n-1}\to\mathbb{R}\mbox{ is Lipschitz-continuous}\},$
endowed with the norm
$\|f\|_{Lip(B_{\rho}^{n-1})}:=\|f\|_{\mathcal{C}^{0}(B_{\rho}^{n-1})}+[f]_{\mathcal{C}^{0}(B_{\rho}^{n-1})},$
@@
$f]_{\mathcal{C}^{0}(B_{\rho}^{n-1})}:=\sup_{\begin{subarray}{c}x^{\prime},y^{\prime}\in B_{\rho}^{n-1}\\ x^{\prime}\neq y^{\prime}\end{subarray}}\frac{|f(x^{\prime})-f(y^{\prime})|}{|x-y^{\prime}|}$
denotes the usual Lipschitz seminorm.
Then, we recall the following notion of Lipschitz set (see [MR3791463, Definition 2.4.5]).
@@
We say that an open set $\Omega\subset R^{n}$ has Lipschitz boundary (or that $\Omega$ is a Lipschitz set) if there exist $r_{\Omega}>0$, $L_{\Omega}>0$, and $\alpha_{\Omega}>0$ such that for every $x_{0}\in\partial\Omega$ there exist a rotation $\mathcal{R}_{\theta_{x_{0}}}$ (of an angle $\theta_{x_{0}}\in[0,2\pi)$) and a Lipschitz function $\phi_{x_{0}}\in\mbox{Lip}(B_{r_{\Omega}}^{n-1})$ such that:
@@
    $\mathcal{R}_{\theta_{x_{0}}}\Omega(-x_{0})\cap\left(B_{r_{\Omega}}^{n-1}\times(-\alpha_{\Omega},\alpha_{\Omega})\right)=\left\{(x^{\prime},x_{n})\mbox{ s.t. }x^{\prime}\in B_{r_{\Omega}}^{n-1}\mbox{ and }x_{n}\in(-\alpha_{\Omega},\phi_{x_{0}}(x^{\prime}))\right\}$;
  3. iii)
    $\partial\left(\mathcal{R}_{\theta_{x_{0}}}\Omega(-x_{0})\right)\cap\left(B_{r_{\Omega}}^{n-1}\times(-\alpha_{\Omega},\alpha_{\Omega})\right)=\left\{(x^{\prime},\phi_{x_{0}}(x^{\prime}))\mbox{ s.t. }x^{\prime}\in B_{r_{\Omega}}^{n-1}\right\}$.
@@
There exist positive constants $\eta\in(0,1)$ and $C=C(n,s_{1},s_{2},\kappa_{2},\delta,\eta,\Omega)$ such that
@@
Another important tool for proving Proposition C.2 is an adaptation of [MR2782803, Lemma 10] for Lipschitz domains. For this, let us define for any $\rho>0$ and $\lambda\geqslant 1$
@@
Let us consider $s\in(0,1)$, $\rho>0$, and $\lambda\geqslant 1$, and let $\phi\in\mbox{Lip}(B_{\rho}^{n-1})$ be a Lipschitz function such that $\phi=0$ and $\nabla\phi(0)=0$.
Let also
and $C_{\rho,\lambda}^{-}(\phi):=\{(x^{\prime},x_{n})\in\mathbb{R}^{n}\mbox{ s.t. }|x^{\prime}|<\rho,\ -\rho\lambda<x_{n}<\phi(x^{\prime})\}).$
@@
then there exist a transformation $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $\Psi(C_{\rho}^{\pm}(\phi))=C_{\rho}^{\pm}$ and
(C.3) $\iint_{C_{\rho,\lambda}^{+}(\phi)\times C_{\rho,\lambda}^{-}(\phi)}\frac{dxdy}{|x-y|^{n+2s}}\leqslant c\iint_{C_{\rho,\lambda}^{+}\times C_{\rho,\lambda}^{-}}\frac{dxdy}{|x-y|^{n+2s}},$
@@
$\sup_{|x^{\prime}|<\rho}\phi(x^{\prime})\leqslant[\phi]_{Lip(B_{\rho}^{n-1})}\rho.$
Then, we define function $\psi\in\mathcal{C}^{\infty}(\mathbb{R}^{n})$ such that
$\psi(x):=\begin{cases}1,\quad&\mbox{if }|x_{n}-\phi(x^{\prime})|\leqslant\frac{\lambda-[\phi]_{Lip(B_{\rho}^{n-1})}}{8}\rho,\\ 0,\quad&\mbox{if }|x_{n}-\phi(x^{\prime})|\geqslant(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho,\end{cases}$
@@
$|\nabla\psi|\leqslant\frac{2}{(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho}.$
Notice that, for any $|x^{\prime}|\leqslant\rho$, we have that
$|\phi(x^{\prime})\pm\lambda\rho|\geqslant\lambda\rho-|\phi(\rho)|\geqslant(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho,$
@@
By Rademacher’s Theorem, $\phi$ is differentiable almost everywhere, and, if $\widetilde{x}:=\Psi(x)$ and $\widetilde{y}:=\Psi(y)$, we have
$\begin{split}&d\widetilde{x}\,d\widetilde{y}=|1-\partial_{n}\Psi(x)\phi(x^{\prime})|\,|1-\partial_{n}\Psi(y)\phi(y^{\prime})|dx\,dy\\ &\qquad\geqslant\left(1-\left|\partial_{n}\Psi(x)\phi(x^{\prime})+\partial_{n}\Psi(y)\phi(y^{\prime})-\partial_{n}\Psi(x)\phi(x^{\prime})\partial_{n}\Psi(y)\phi(y^{\prime})\right|\right)dx\,dy\\ &\qquad\geqslant(1-c_{1})dx\,dy,\end{split}$
(see also [MR3409135, Theorem 2, Sec. 3.3.3]) with
@@
where we have employed (C.2) to obtain the last inequality. It thereby follows that
(C.7) $dx\,dy\leqslant\frac{1}{1-c_{1}}d\widetilde{x}\,d\widetilde{y}$.

### v2

@@ 9. Proof of the $\Gamma$-$\liminf$ inequality
$\mbox{and}\quad\lambda(A)\leqslant\liminf_{\varepsilon\to 0}\lambda_{\varepsilon}(A),\quad\mbox{for every set\penalty 10000\ $A$ such that\penalty 10000\ $\lambda(\partial A)=0$.}$
@@
Now, let $\overline{x}$ be a regular point of $E$, namely a point $\overline{x}\in\partial E$ with (n1)(n-1)-density 11 and such that the blow-up of $E$ around $\overline{x}$ converges to $\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}$. More precisely,
@@
and $\lim_{r\to 0^{+}}\frac{1}{r^{n}}\int_{B_{2r}(\overline{x})}|\chi_{E}-\chi_{\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}}(x)|\,dx=0,$
where $\omega_{n-1}:=\mathcal{H}^{n-1}(B^{n-1}_{1})$.
We consider the Radon-Nikodyn derivative of $\lambda$ with respect to the (n1)(n-1)-dimensional Hausdorff measure at $\overline{x}$, namely
@@
(9.3) $\frac{d\lambda}{d\mathcal{H}^{n-1}}(\overline{x})\geqslant\phi(\nu_{E}(\overline{x})),$
where $\phi$ is as in (1.21).
Let us consider subsequences $\{\varepsilon_{j}\}_{j}$, $\{r_{j}\}_{j}$, and $\{\eta_{j}:=\varepsilon_{j}/r_{j}\}_{j}$ such that
(9.4) $\begin{split}&\lim_{j\to+\infty}\frac{\lambda_{\varepsilon_{j}}(B_{r_{j}}(\overline{x}))}{\omega_{n-1}r_{j}^{n-1}}=\frac{d\lambda}{d\mathcal{H}^{n-1}}(\overline{x}),\\ &\lim_{j\to+\infty}\frac{1}{r_{j}^{n}}\int_{B_{2r_{j}}(\overline{x})}|\chi_{E}(x)-\chi_{\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}}(x)|\,dx=0,\\ \mbox{and}\quad&\lim_{j\to+\infty}\eta_{j}=0.\end{split}$
In what follows, in order to lighten the notation, we will write $E_{j}:=E_{\varepsilon_{j}}$, $\lambda_{\varepsilon_{j}}:=\lambda_{j}$, and $\nu:=\nu_{E}(\overline{x})$. We now focus on proving (9.3), that in this notation reads
@@
$\begin{split}\frac{\lambda_{j}(B_{r_{j}}(\overline{x}))}{r_{j}^{n-1}}&=\iint_{B_{1}(\eta_{j}\{\overline{x}/\varepsilon_{j}\})\times\mathbb{R}^{n}}\chi_{E^{\prime\prime}_{j}}(x)\chi_{E^{\prime\prime c}_{j}}(y)K_{\eta_{j}}(x,y)\,dx\,dy+\int_{E^{\prime\prime}_{j}\cap\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}}g_{\eta_{j}}(x)\,dx\\ &=\iint_{B^{\prime\prime}_{j}\times\mathbb{R}^{n}}\chi_{E^{\prime\prime}_{j}}(x)\chi_{E^{\prime\prime c}_{j}}(y)K_{\eta_{j}}(x,y)\,dx\,dy+\int_{E^{\prime\prime}_{j}\cap\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}}g_{\eta_{j}}(x)\,dx,\end{split}$
where, to be consistent with our notation, we wrote $B_{j}^{\prime\prime}:=B_{1}(\eta_{j}\{\overline{x}/\varepsilon_{j}\})$.
Moreover, observe that
$\begin{split}\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}&=\bigcup\left\{\frac{\varepsilon_{j}(k+Q)-\overline{x}}{r_{j}}+\eta_{j}\left\{\frac{\overline{x}}{\varepsilon_{j}}\right\}\mbox{s.t. }k\in\mathbb{Z}^{n},\ \varepsilon_{j}(k+Q)\subseteq B_{r_{j}}(\overline{x})\right\}\\ &=\bigcup\left\{\eta_{j}\left(k-\left\lfloor\frac{\overline{x}}{\varepsilon_{j}}\right\rfloor+Q\right)\mbox{s.t. }k\in\mathbb{Z}^{n},\ \eta_{j}\left(k-\left\lfloor\frac{\overline{x}}{\varepsilon_{j}}\right\rfloor+Q\right)\subseteq B^{\prime\prime}_{j}(\overline{x})\right\}\\ &=\mathcal{Q}(B^{\prime\prime}_{j})_{\eta_{j}}.\end{split}$
@@
(9.7) $\begin{split}&\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{E^{\prime\prime}_{j}}(x)-\chi_{\mathcal{I}_{\nu}}(x)|\,dx\\ \leqslant\;&\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{E^{\prime\prime}_{j}}(x)-\chi_{\{(x-\eta_{j}\{\overline{x}/\varepsilon_{j}\})\cdot\nu>0\}}(x)|\,dx\\ &\qquad+\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{\{(x-\eta_{j}\{\overline{x}/\varepsilon_{j}\})\cdot\nu>0\}}(x)-\chi_{\mathcal{I}_{\nu}}(x)|\,dx\\ =\;&0.\end{split}$
@@
(9.8) $F_{j}:=(\eta_{j}E_{\nu}\setminus\mathcal{Q}(B^{\prime\prime}_{j})_{1})\cup(E^{\prime\prime}_{j}\cap\mathcal{Q}(B^{\prime\prime}_{j})_{1}),$
see Figure 2.
@@
(9.9) $\lim_{j\to+\infty}\mathscr{F}_{\eta_{j}}(F_{j},B^{\prime\prime}_{j})\geqslant\lim_{j\to+\infty}\mathscr{F}_{\eta_{j}}(\eta_{j}E_{\nu},B^{\prime\prime}_{j})=\mathcal{H}^{n-1}(\partial\mathcal{I}_{\nu}\cap B_{1})\phi(\nu)=\omega_{n-1}\phi(\nu).$
@@ 10. Continuity of the stable norm $\phi$
$\begin{split}\varepsilon^{1-n}\zeta(\varepsilon):=&\left|S^{q}_{M+\sqrt{n}}\cap S^{p}_{M+\sqrt{n}}\cap\mathcal{R}_{\theta}E^{q}\cap\big(C^{p}/\varepsilon\setminus\mathcal{Q}(Q^{p}/\varepsilon)_{1}\big)\right|\\ &+\left|S^{q}_{M+\sqrt{n}}\cap S^{p}_{M+\sqrt{n}}\cap E^{q}\cap\big(C^{q}/\varepsilon\setminus\mathcal{Q}(Q^{q}/\varepsilon)_{1}\big)\right|,\end{split}$
@@
$\begin{split}&\phi(q)-\phi(p)\\ \geqslant&\liminf_{\varepsilon\to 0}\left(-\varepsilon^{n-1}\|g\|_{L^{\infty}(\mathbb{R}^{n})}\left|\left(E_{q}\cap S^{q}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{q}\right)\Delta\left(\mathcal{R}_{\theta}E_{q}\cap S^{p}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{p}\right)\right|\right).\end{split}$
@@
$\psi(\theta):=\limsup_{\varepsilon\to 0}\varepsilon^{n-1}\left|\left(E_{q}\cap S^{q}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{q}\right)\Delta\left(\mathcal{R}_{\theta}E_{q}\cap S^{p}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{p}\right)\right|,$
@@
$\phi(q)-\phi(p)\geqslant-\|g\|_{L^{\infty}(\mathbb{R}^{n})}\psi(\theta).$
@@
Interchanging the roles of $p$ and $q$, we obtain that
$\begin{split}&\phi(p)-\phi(q)\geqslant\liminf_{\varepsilon\to 0}\left(-\|g\|_{L^{\infty}(\mathbb{R}^{n})}\left|\big(E_{q}\cap S^{p}_{M+\sqrt{n}}\cap C^{p}\big)\Delta\big(\mathcal{R}_{-\theta}\cap S^{q}_{M+\sqrt{n}}\cap C^{q}\big)\right|\right)\\ &\qquad\qquad=-\|g\|_{L^{\infty}(\mathbb{R}^{n})}|J\mathcal{R}_{\theta}|\psi(\theta)=-\|g\|_{L^{\infty}(\mathbb{R}^{n})}\psi(\theta),\end{split}$
@@ 11. Proof of the $\Gamma$-$\limsup$ inequality
We point out that, in light of Lemma 7.1, one can equivalently consider energies $\mathscr{F}_{\varepsilon}((\cdot),\Omega)$, $\mathscr{E}_{\varepsilon}((\cdot),\Omega)$, or even $\mathscr{L}_{\varepsilon}((\cdot),\Omega):=\mathcal{L}_{K_{\varepsilon}}((\cdot)\cap\Omega,(\cdot)^{c}\cap\Omega)$ in Theorem 1.13. Indeed, since the rescaled planelike $\varepsilon E_{p}$ converges to the straight line $\mathcal{I}_{p}$, as $\varepsilon\to 0$, such limit crosses transversely any Lipschitz domain. Therefore, the energy contribution due to the interactions with the exterior of $\Omega$ vanishes in the limit.
@@ Appendix A Norms subject to short-range interactions
We point out a useful equivalence of norms when a kernel is restricted to close-by interactions (see e.g. [2025arXiv250409976A, Lemma 2.1] for related results). For the rest of this section, let $s\in(0,1/2)$.
@@ Appendix C Behavior of $\operatorname{P}_{K}$ in large domains
First, let us consider the space of Lipschitz function in $B_{\rho}^{n-1}$, for some $\rho>0$, denoted by $\mbox{Lip}(B_{\rho}^{n-1})$ and endowed with the norm
$\|f\|_{Lip(B_{\rho}^{n-1})}:=\|f\|_{L^{\infty}(B_{\rho}^{n-1})}+[f]_{Lip(B_{\rho}^{n-1})},$
@@
$f]_{Lip(B_{\rho}^{n-1})}:=\sup_{\begin{subarray}{c}x^{\prime},y^{\prime}\in B_{\rho}^{n-1}\\ x^{\prime}\neq y^{\prime}\end{subarray}}\frac{|f(x^{\prime})-f(y^{\prime})|}{|x-y^{\prime}|}.$
We also recall the following notion of Lipschitz set (see [MR3791463, Definition 2.4.5]).
@@
We say that an open set $\Omega\subset\mathbb{R}^{n}$ has Lipschitz boundary (or that $\Omega$ is a Lipschitz set) if there exist $r_{\Omega}>0$, $L_{\Omega}>0$, and $\alpha_{\Omega}>0$ such that for every $x_{0}\in\partial\Omega$ there exist a rotation $\mathcal{R}_{\theta_{x_{0}}}$ (of an angle $\theta_{x_{0}}\in[0,2\pi)$) and a Lipschitz function $\phi_{x_{0}}\in\mbox{Lip}(B_{r_{\Omega}}^{n-1})$ such that:
@@
    $\mathcal{R}_{\theta_{x_{0}}}(\Omega-x_{0})\cap\left(B_{r_{\Omega}}^{n-1}\times(-\alpha_{\Omega},\alpha_{\Omega})\right)=\left\{(x^{\prime},x_{n})\mbox{ s.t. }x^{\prime}\in B_{r_{\Omega}}^{n-1}\mbox{ and }x_{n}\in(-\alpha_{\Omega},\phi_{x_{0}}(x^{\prime}))\right\}$.
@@
There exist $\eta\in(0,1)$ and $C=C(n,s_{1},s_{2},\kappa_{2},\delta,\eta,\Omega)>0$ such that
@@
Another important tool for proving Proposition C.2 is an adaptation of [MR2782803, Lemma 10] for Lipschitz domains. For this, let us define, for any $\rho>0$ and $\lambda\geqslant 1$, the sets
@@
For any $\phi\in\mbox{Lip}(B_{\rho}^{n-1})$, let also
and $C_{\rho,\lambda}^{-}(\phi):=\{(x^{\prime},x_{n})\in\mathbb{R}^{n}\mbox{ s.t. }|x^{\prime}|<\rho,\ -\rho\lambda<x_{n}<\phi(x^{\prime})\}.$
@@
Let $s\in(0,1)$, $\rho>0$, and $\lambda\geqslant 1$, and let $\phi\in\mbox{Lip}(B_{\rho}^{n-1})$ be such that $\phi(0)=0$ and $\nabla\phi(0)=0$.
@@
then there exists a transformation $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $\Psi(C_{\rho}^{\pm}(\phi))=C_{\rho}^{\pm}$ and
$\iint_{C_{\rho,\lambda}^{+}(\phi)\times C_{\rho,\lambda}^{-}(\phi)}\frac{dx\,dy}{|x-y|^{n+2s}}\leqslant c\iint_{C_{\rho,\lambda}^{+}\times C_{\rho,\lambda}^{-}}\frac{dx\,dy}{|x-y|^{n+2s}},$
@@
(C.3) $\sup_{|x^{\prime}|<\rho}|\phi(x^{\prime})|\leqslant[\phi]_{Lip(B_{\rho}^{n-1})}\rho.$
We recall (C.2) and we define the function $\psi\in\mathcal{C}^{\infty}(\mathbb{R}^{n},[0,1])$ such that
$\psi(x):=\begin{cases}1,\quad&\mbox{if }\displaystyle|x_{n}-\phi(x^{\prime})|\leqslant\frac{\lambda-[\phi]_{Lip(B_{\rho}^{n-1})}}{8}\rho,\\ 0,\quad&\mbox{if }\displaystyle|x_{n}-\phi(x^{\prime})|\geqslant\big(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})}\big)\rho,\end{cases}$
@@
(C.4) $|\nabla\psi|\leqslant\frac{2}{(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho}.$
Notice that, for any $|x^{\prime}|\leqslant\rho$, by (C.3) we have that
$|\phi(x^{\prime})\pm\lambda\rho|\geqslant\lambda\rho-|\phi(x^{\prime})|\geqslant(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho$
@@
By Rademacher’s Theorem, $\phi$ is differentiable almost everywhere, and, if $\widetilde{x}:=\Psi(x)$ and $\widetilde{y}:=\Psi(y)$, we have (see also [MR3409135, Theorem 2, Sec. 3.3.3])
$\begin{split}d\widetilde{x}\,d\widetilde{y}&=|1-\partial_{n}\psi(x)\phi(x^{\prime})|\,|1-\partial_{n}\psi(y)\phi(y^{\prime})|dx\,dy\\ &\geqslant\left(1-\left|\partial_{n}\psi(x)\phi(x^{\prime})+\partial_{n}\psi(y)\phi(y^{\prime})-\partial_{n}\psi(x)\phi(x^{\prime})\partial_{n}\psi(y)\phi(y^{\prime})\right|\right)dx\,dy\\ &\geqslant(1-c_{1})dx\,dy,\end{split}$
with
@@
where we have employed (C.2), (C.3) and (C.4). It thereby follows that
(C.8) $dx\,dy\leqslant\frac{d\widetilde{x}\,d\widetilde{y}}{1-c_{1}}.$